# Making Every Frame Matter: Continuous Video Understanding for Large Models via Adaptive State Modeling

### v1

@@ 5.2. Evaluation of All Five Capabilities
We chose the most natural input strategy for each baseline to assess the accuracy. For batch-based methods, i.e., Baselines 1, 2, 4, and 5, we adopted the all-frame feeding strategy. Although this significantly increases computational overhead, it enables these methods to have sufficient context to understand the video content. Both Baseline 3 and our method utilize a streaming processing paradigm, for which we selected the single-frame feeding strategy.
In the evaluation of single-label classification datasets, we utilize Accuracy (Acc) as the primary measure of correctness. This accuracy is ascertained by verifying whether the correct action is detected within a given video. The EDR is calculated as introduced in Section 5.1(Metrics). When dealing with multi-label classification datasets, two variants of mean Average Precision (mAP) are employed to gauge correctness. The $\text{mAP}_{\text{all}}$ metric is determined by considering predictions of all frames in the video. The $\text{mAP}_{\text{last}}$ entails calculations based solely on the prediction results of the final frame in the video. It’s important to note that the label of the last frame includes labels for all actions present within the current video. Each label’s EDR is computed separately and then averaged. If a method fails to detect a particular label, this failure is taken into account during computation by adding a 100% increase to the result for that category. This signifies that the correct label was not detected even upon the conclusion of the video.
Below we present our experimental analyses and the subsequent conclusions:
Superior Accuracy: Our method consistently outperforms all baseline methods in terms of accuracy across all tasks, improving metrics by ranges of 2 to 79.7, 2.9 to 78.2, 1.2 to 14.5, and 0.7 to 16.3. The leading accuracy demonstrates that our C-VUE can effectively understand events of different scales (Capability 1). The superior accuracy on the FineGym288 dataset underscores our method’s generalization ability for long-tailed data (Capability 3)).
Prompt Event Detection: Our approach exhibits an outstanding ability to promptly detect events (Capability 2) and continuously processing capability (Capability 4) by EDR. For single-label tasks, events are detected within the first quarter of their occurrence. For complex multi-label tasks, all events are identified halfway through the video.
Memorization Capability: Within the multi-label dataset Charades, the $\text{mAP}_{\text{last}}$ denotes the ability to recognize all events in the video by its conclusion. Notably, Baselines 1 and 2 attained such accuracy by taking as input all video frames simultaneously. Our method achieves a higher accuracy based on the last frame and implicit memorization of past frames. Notably, our approach achieves 1.24 times the accuracy of Baseline 3, which employs LSTM for continuous processing. Results show that our C-VUE shows a better long-range memorization ability (Capability 5).
@@
Robust Stability Across Varied Tasks: Our method showcases robust stability in tackling diverse tasks. It performs uniformly well across differing task scales (including those with varied event durations), varied video characteristics (motion-focused and scene-focused video), and differing classification objectives (such as single-label and multi-label classifications). No baseline methods are currently able to match this level of comprehensive performance.
@@
Table 2. Evaluation results on All-event-in-one FineGym. (The parenthesized number of S2 signifies the window size, while 725 refers to the maximum frame count of a single event.)
@@
5.3. Further Evaluation of Capabilities 1 & 4
We assess the continuous perception capability with the All-event-in-one FineGym dataset, a 3.5-hour video described in Section 5.1. All baseline methods use the model trained on the FineGym99 dataset, whose accuracy evaluated on a single video clip is reported in Table 2. These methods are evaluated with different input strategies. For batch processing approaches (Baselines 1 and 2), we test all three strategies. Regarding streaming processing methods (our C-VUE and Baseline 3), we use the single frame feeding strategy (Strategy 3). Both our C-VUE and Baseline 3 benefit from the model’s memorization of past video frames. During continuous perception, each frame is processed only once without redundant input. The Acc metric is calculated by considering predictions of all frames in the video.
@@ 5.3. Further Evaluation of Capabilities 1 & 4
Based on the results, we draw the following two conclusions. First, our C-VUE demonstrates the best continuous perception ability (Capability 4). It surpasses batch processing methods ranging from 0.5 to 10.7. Its accuracy is twice that of traditional LSTM-based baseline methods. Second, our approach can effectively adapt to events of varying scales (Capability 1). In the hours-long video, events lasted from tens of frames to hundreds of frames. Our C-VUE achieves the best accuracy, which outperforms the current sliding-window and state-assist methods.
5.4. Further Evaluation of Capability 5
We further profile the memorization abilities on three multi-label Finegym datasets to evaluate our memorization capability. The k value represents how many events are concatenated to form a testing data sample. Therefore, a higher k value indicates a longer video with more labels.
All baseline methods use the model trained on the FineGym99 dataset. Baselines 1 and 2 are evaluated with input strategy 1, where they can achieve the best accuracy. We utilize three metrics. The Jaccard Index [^1] The threshold ‘th’ signifies that in the model’s output vector, values greater than ‘th’ are considered as detecting a particular category, otherwise not recognized. is calculated as the intersection of predicted results and ground truth divided by their union. We report the results in Table 3 and draw the following conclusions.
Our C-VUE demonstrates an superior memorization capabilities (Capability 5). Across all configurations and metrics, our method surpasses the baseline approaches. Our C-VUE excels at tasks involving frame-by-frame understanding. Surprisingly, baseline methods that process frames in batches (Baselines 1 and 2), incorporating all frames at each time step, exhibit notably lower accuracy compared to streaming methods (Ours and Baseline 3). This difference arises from the fact that streaming methods prioritize the present frame while retaining information from past frames in streaming data processing.
Table 3. Evaluation results on three Multi-label FineGym@k datasets.
Metrics Methods k=1 k=3 k=5
Jaccard Index (th=0.5) Baseline 1 0.0351 0.0172 0.0087
Baseline 2 0.0194 0.0281 0.0313
Baseline 3 0.0906 0.0484 0.0287
Ours 0.1210 0.0578 0.0341
Jaccard Index (th=0.3) Baseline 1 0.0803 0.04396 0.0228
Baseline 2 0.0101 0.0152 0.0200
Baseline 3 0.1426 0.0880 0.0554
Ours 0.1752 0.0957 0.0604
Jaccard Index (th=0.1) Baseline 1 0.1324 0.0802 0.0441
Baseline 2 0.0101 0.0152 0.0200
Baseline 3 0.1428 0.1254 0.0935
Ours 0.1846 0.1334 0.0936
mAP Baseline 1 0.1590 0.1196 0.0905
Baseline 2 0.0465 0.0374 0.0353
Baseline 3 0.2338 0.1645 0.1277
Ours 0.2678 0.1739 0.1379
ranking AP Baseline 1 0.3879 0.3296 0.2771
Baseline 2 0.4533 0.4338 0.4108
Baseline 3 0.5856 0.4906 0.4250
Ours 0.5894 0.4916 0.4346
Figure 9. Processing speed of understanding a video with 10k frames on different platforms.
5.5. Real-time Inference on Edge
@@
We measure the system performance on three mobile platforms with two long video clips. The long video clips were excerpted from the All-event-in-one FineGym dataset, one containing 10k frames and the other 100k frames. We conducted each experiment three times and calculated the average. In Figure 9, we present the system performance of different baselines processing 10k frame videos on the three platforms. Additionally, in Figure 4, we report the system performance of different baselines processing 100k frame videos on a PC equipped with an RTX4090. FPS is calculated by dividing the total number of video frames by the total processing time. Our experimental findings are as follows.
@@
Understanding the video stream in real time. In Figure 9, it can be observed that our method achieves real-time processing (¿30fps) on different platforms. Even on Macbook Air and Jetson, our speed is 2 to 20 times faster than the baseline methods. This improvement stems from two aspects. Firstly, our streaming processing method eliminates the need to use the model through all feeding or sliding window approaches. Secondly, in terms of system implementation and optimization, recall that mainstream frameworks, e.g., OpenClip, clip.cpp, and ONNXRuntime-GPU, lack support for using the CLIP model with Jetson GPU. Therefore, baselines can only utilize the CPU for computation. Our prototype can enable GPU computation on both Macbook Air and Jetson.
@@
6. Case Study: Video Encoder for LLM-based Video Foundation Models
Figure 10. Applying our history modeling into VideoLLaMa2 as a video encoder.
In this section, we discuss how the design of our system, C-VUE, improves video foundation models. As a case study, we analyze one of the SOTA video foundation models, VideoLLaMa2 (damonlpsg2024videollama2, ) (left side of Figure 10). During video processing, VideoLLaMa2 first samples $N$ frames from the video. These frames are then converted into $M$ tokens by a multimodal projector (denoted as ”Proj.” in the figure). Simultaneously, the user’s text prompt is tokenized, combined with the video tokens, and input into the LLM for inference.
6.1. Problem Analysis
@@
6.2. Our design
We adopt the design of our C-VUE to build a new projection layer, named the C-VUE projector, which replaces the STC projector. This new projection utilizes the long-range history modeling to enable a high sampling rate in a remarkably efficient manner. Figure 10 (right) illustrates our design.
When a video frame is input, the projector first processes the spatial information within the frame by applying mean pooling to patch embeddings. Next, the history modeling layer processes the history information, capturing relationships over time. Its processing speed remains consistent regardless of video length, making it more efficient.
Table 4. Evaluation results on the VideoLLaMa2.
@@ 6.3. Experiments
Implementation. Using the official videollama2 codebase (video_llama_impl, ), we implemented our C-VUE projector. This projector consists of mean pooling, a linear layer with ReLU, and the history modeling layer of our C-VUE with a code dimension of 2048, followed by another linear layer with ReLU. The total parameter count of our projector is 450 M. For comparison, the STC projector with 8 frames has 1.8 B parameters, which is four times larger.
@@
We fine-tuned and evaluated the model following the official guidelines. All hyperparameters were set to the official default values. During fine-tuning, only the large language model (LLM) and the projection layer were trained. We used the videoChatGPT dataset (maaz2023video, ) for fine-tuning both the original model structure and our proposed model structure. All experiments were performed on a machine with eight A600 GPUs. Due to limited training resources, we used the 8-frame model as the baseline, while our method extended the training sample length to 128 frames.
@@
Results. Testing was conducted on the videoChatGPT dataset (in-distribution), ActivityNet-QA (yu2019activityqa, ) (zero-shot), and MSVD-QA (xu2017video, ) (zero-shot), and the results are reported in Table 4. For the VideoChatGPT dataset, correctness was measured on a 5-point scale. Our method outperformed the baseline by 0.46 points. On the other two zero-shot datasets, our approach achieved improvements of 1.19% and 4%, respectively. These results demonstrate that our projector design effectively improves model accuracy, highlighting a new design concept for video projection for video foundation models.
7. Related Work
Video Analytics & Understanding. Video analytics and understanding involve the automatic extraction of valuable information from video including classification (wu2020emo, ), detection (tong2021large, ; jiang2021flexible, ), tracking (voigtlaender2019mots, ), identification (yi2020eagleeye, ), and privacy protection (wu2021pecam, ). Many approaches treat continuous video analysis as a task of image understanding on video frames. Adaptive frame sampling (hwang2022cova, ; moll2022exsample, ), cooperation between small and large models (liu2020continuous, ), and collaboration across multiple computation resources (chen2015glimpse, ; liu2019edge, ; damonlpsg2024videollama2, ) are proposed to enable efficient video analytics on video frames.
To analyze the events that involve multiple frames, 3D convolution is used to perform the video understanding (slowfast2019fei, ). However, this approach can cause significant system performance issues (zhu2020comprehensive, ). To improve efficiency, MoviNet (kondratyuk2021movinets, ) creates efficient 3D convolution architectures. Many efforts also explore temporal modeling without 3D convolution (Lin_2019_ICCV, ; Jiang_2019_ICCV, ; Li_2020_CVPR, ). MViTv2 (li2022mvitv2, ) uses improved pooling attention for spatiotemporal recognition to improve the attention’s efficiency. These methods improve video understanding by focusing on relationships in a short video clip. They still struggle with system performance issues in long-form and streaming videos.
Video Foundation Models. Recent studies have explored building video foundation models using large language models (LLMs). These models achieve impressive results in multiple video understanding tasks (madan2024foundation, ).
Methods like VideoLaVIT (jin2024video, ), VideoLLaVA (lin2023video, ), Video-LLaMA (zhang2023video, ), and Video-LLaMA2 (damonlpsg2024videollama2, ) use multi-modal pre-training to align visual information with LLMs through projection layers. VideoChatGPT (maaz2023video, ) improves video-based conversations by integrating visual encoders with LLMs and instruction tuning. The main challenge in video understanding is handling memory requirements and model capacity, especially for long videos. LaViLa (zhao2023learning, ) addresses this by turning LLMs into narrators, allowing them to generate detailed descriptions of long videos from visual inputs. ChatVideo (wang2023chatvideo, ) and MM-VID (lin2023mm, ) convert videos into text for better comprehension. However, these methods still struggle to model long-range temporal information in streaming videos.
Alternatives to Foundation Models for Video Understanding. The series of works on Mamba have gained significant attention for overcoming the high computational complexity of Transformers as LLMs. Recently, Mamba has been applied to various computer vision tasks, showing early progress (liu2024vision, ). However, its application is still in its infancy. Mamba-ND (li2024mambandselectivestatespace, ) is the first to demonstrate Mamba can handle high-dimensional data and experimentally show the effectiveness in video action recognition. VideoMamba (li2024videomambastatespacemodel, ) builds an SSM-based model for video understanding. It shares a similar design with VideoLlama (damonlpsg2024videollama2, ) but replaces the Transformer-based LLM with a Mamba-based one for efficiency purposes. VideoMamba Suite (chen2024video, ) further evaluates Mamba’s performance by integrating it into existing video understanding architectures, showing its potential to handle complex spatial-temporal dynamics.
Unlike these works that use “all frames at once” or “fixed window size” strategies for Mamba and Transformer video understanding, we propose a lightweight continuous video understanding design with each frame as the input to the state-assistance module.
8. Discussion & Conclusion
Discussion. Our work presents a solution for continuous video understanding, but there are still unexplored issues we hope to address with the community in the future. The memorization upper bound of the state modeling approach has not been fully explored. Although several improvements were made, a fixed-size hidden state will fail to model infinitely long videos. Future work should figure out the modeling capacity of such methods. Additionally, the generalization ability of these approaches to more and more tasks remains unclear, which is crucial for designing new video foundation models based on this design.
@@ 8. Discussion & Conclusion
Conclusion. Continuous video understanding is important for complex multimodal human-computer interaction. However, achieving this on resource-limited edge devices is challenging. This paper proposes C-VUE, an adaptive state modeling approach, for continuous video understanding. Experimental results show that C-VUE can achieve real-time continuous video understanding on typical edge platforms. As a case study, we apply C-VUE’s design to improve video foundation models and obtain a large accuracy improvement.

### v2

@@ 4.2 Evaluation of All Capabilities
We chose the most natural input strategy for each baseline to assess the accuracy. For batch-based methods, i.e., Baselines 1, 2, 4, and 5, we adopted the all-frame feeding strategy. Although this significantly increases computational overhead, it enables these methods to have sufficient context to recognize the video activities. Both Baseline 3 and our method utilize a streaming processing paradigm, for which we selected the single-frame feeding strategy.
In evaluating single-label classification datasets, we utilize Accuracy (Acc) as the primary measure of correctness. This accuracy is ascertained by verifying whether the correct action is detected within a given video. The EDR is calculated as introduced in Section 4.1(Metrics). When dealing with multi-label classification datasets, two variants of mean Average Precision (mAP) are employed to gauge correctness. The $\text{mAP}_{\text{all}}$ metric is determined by considering predictions of all frames in the video. The $\text{mAP}_{\text{last}}$ entails calculations based solely on the prediction results of the final frame in the video. It’s important to note that the label of the last frame includes labels for all actions present within the current video. Each label’s EDR is computed separately and then averaged. If a method fails to detect a particular label, this failure is taken into account during computation by adding a 100% increase to the result for that category. This signifies that the correct label was undetected even after the video’s conclusion. Below we present our experimental analyses and the subsequent conclusions:
Superior Accuracy: Our method consistently outperforms all baseline methods in terms of accuracy across all tasks, improving metrics by ranges of 2 to 79.7, 2.9 to 78.2, 1.2 to 14.5, and 0.7 to 16.3. The leading accuracy demonstrates that our CARS can effectively recognize activities of different scales (Capability 1). The superior accuracy on the FineGym288 dataset underscores our method’s generalization ability for long-tailed data (Capability 4)).
Prompt Recognition : Our approach exhibits an outstanding ability to promptly detect activities (Capability 3) and continuously process (Capability 2) by EDR. For single-label tasks, activities are detected within the first quarter of their occurrence. For complex multi-label tasks, all activities are identified halfway through the video.
@@
Robust Stability Across Varied Tasks: Our method showcases robust stability in tackling diverse tasks. It performs uniformly well across differing task scales (including those with varied activity durations), varied video characteristics (motion-focused and scene-focused video), and differing classification objectives (such as single-label and multi-label classifications). No baseline methods are currently able to match this level of comprehensive performance.
Maintaining Video Context for Multiple Activities: Within the multi-label dataset Charades, the $\text{mAP}_{\text{last}}$ denotes the ability to recognize all activities in the video by its conclusion. Notably, Baselines 1 and 2 attained such accuracy by taking as input all video frames simultaneously. Our method achieves a higher accuracy based on the last frame and implicit memorization of past frames. Notably, our approach achieves 1.24 times the accuracy of Baseline 3, which employs LSTM for continuous processing. Results show that our CARS shows a better utilization of the hidden state.
@@
Table 2: Evaluation results on All-activity-in-one FineGym. (The parenthesized number of S2 signifies the window size, while 725 refers to the maximum frame count of a single activity.)
@@
Figure 9: Processing speed of streaming video activity recognition with 10k frames on different platforms.
4.3 Further Evaluation of Capabilities 1 & 2
We assess the continuous perception capability with the All-activity-in-one FineGym dataset, a 3.5-hour video described in Section 4.1. All baseline methods use the model trained on the FineGym99 dataset, whose accuracy evaluated on a single video clip is reported in Table 2. These methods are evaluated with different input strategies. For batch processing approaches (Baselines 1 and 2), we test all three strategies. Regarding streaming processing methods (our CARS and Baseline 3), we use the single frame feeding strategy (Strategy 3). Both our CARS and Baseline 3 benefit from the model’s memorization of past video frames. During continuous perception, each frame is processed only once without redundant input. The Acc metric is calculated by considering predictions of all video frames.
@@ 4.3 Further Evaluation of Capabilities 1 & 2
Based on the results, we draw the following two conclusions. First, our CARS demonstrates the best continuous perception ability (Capability 2). It surpasses batch processing methods ranging from 0.5 to 10.7. Its accuracy is twice that of traditional LSTM-based baseline methods. Second, our approach can effectively adapt to activities of varying scales (Capability 1). In the hours-long video, activities lasted from tens of frames to hundreds of frames. Our CARS achieves the best accuracy, which outperforms the current sliding-window and state-assist methods.
4.4 Real-time Inference on Edge
@@
We measure the system performance on three mobile platforms with two long video clips. The long video clips were excerpted from the All-activity-in-one FineGym dataset, one containing 10k frames and the other 100k frames. We conducted each experiment three times and calculated the average. In Figure 9, we present the system performance of different baselines processing 10k frame videos on the three platforms. Additionally, in Figure 4, we report the system performance of different baselines processing 100k frame videos on a PC equipped with an RTX4090. FPS is calculated by dividing the total number of video frames by the total processing time.
Our experimental findings are as follows.
@@
Recognizing the video stream in real time. In Figure 9, it can be observed that our method achieves real-time processing ($>$30fps) on different platforms. Even on Macbook Air and Jetson, our speed is 2 to 20 times faster than the baseline methods. This improvement stems from two aspects. Firstly, our streaming processing method eliminates the need to use the model through all feeding or sliding window approaches. Secondly, in terms of system implementation and optimization, recall that mainstream frameworks, e.g., OpenClip, clip.cpp, and ONNXRuntime-GPU, lack support for using the CLIP model with Jetson GPU. Therefore, baselines can only utilize the CPU for computation. Our prototype can enable GPU computation on both Macbook Air and Jetson.
@@
5 Case Study: Video Encoder for Large Video Models
Figure 10: Applying our CARS into VideoLLaMa2 as a video encoder.
In this section, we discuss how the design of our system, CARS, improves large video models in video activity recognition and understanding. As a case study, we analyze one of the SOTA large video models, VideoLLaMa2 [8] (left side of Figure 10). During video processing, VideoLLaMa2 first samples $N$ frames from the video. These frames are then converted into $M$ tokens by a multimodal projector (denoted as ”Proj.” in the figure). Simultaneously, the user’s text prompt is tokenized, combined with the video tokens, and input into the LLM for inference.
5.1 Problem Analysis
@@
5.2 Our design
We adopt the design of our CARS to build a new projection layer, named the CARS projector, which replaces the STC projector. This new projection utilizes our activity-aware state update to maintain the video context efficiently. Our state-assisted design can efficiently increase the sampling rate of the input video, far exceeding the existing designs with 8 or 16 frames, without increasing the processing time. Figure 10 (right) illustrates our design. When a video frame is input, the projector first processes the spatial information within the frame by applying mean pooling to patch embeddings. Next, our state-assistant module processes the input video, capturing relationships over time.
Table 3: Evaluation results on the VideoLLaMa2.
@@ 5.3 Experiments
Implementation. Using the official videollama2 codebase [6], we implemented our CARS projector. This projector consists of mean pooling, a linear layer with ReLU, and the history modeling layer of our CARS with a code dimension of 2048, followed by another linear layer with ReLU. The total parameter count of our projector is 450 M. For comparison, the STC projector with 8 frames has 1.8 B parameters, which is four times larger.
@@
We fine-tuned and evaluated the model following the official guidelines. All hyperparameters were set to the official default values. During fine-tuning, only the large language model (LLM) and the projection layer were trained. We used the videoChatGPT dataset [25] for fine-tuning both the original model structure and our proposed model structure. All experiments were performed on a machine with eight A600 GPUs. Due to limited training resources, we used the 8-frame model as the baseline, while our method extended the training sample length to 128 frames.
@@
Results. Testing was conducted on the videoChatGPT dataset (in-distribution), ActivityNet-QA [44] (zero-shot), and MSVD-QA [43] (zero-shot), and the results are reported in Table 3. For the VideoChatGPT dataset, correctness was measured on a 5-point scale. Our method outperformed the baseline by 0.46 points. On the other two zero-shot datasets, our approach achieved improvements of 1.19% and 4%, respectively. These results demonstrate that our projector design effectively improves model accuracy, highlighting a new design concept for video projection for large video models.
6 Conclusion
@@
Continuous video activity recognition in streaming video is important for complex multimodal human-computer interaction, e.g., embodied AI. However, achieving this on resource-limited edge devices is challenging. This paper proposes CARS, a state-assistant approach for continuous video activity recognition in streaming video. It consists of two key designs: activity spatial feature extraction to reduce spatial redundancy and activity-aware state update to maintain the video context effectively. Experimental results show that CARS can achieve real-time continuous video recognition on typical edge platforms. We believe that our CARS offers a novel idea for future streaming video processing.